OpenStudy (anonymous):

When the tension on a wire is kept constant, the number of vibrations per second varies inversely as the length of the wire. A 320 cm wire vibrates 420 times per second. How long should the wire be so that it will vibrate 480 times?

OpenStudy (anonymous):

V = k/l k is constant 420 = k/320 Solve for k k = 420 * 320 =134400 v = 134400/ l V = 480 480 = 134400/ i l = 134400/ 480 = 280 The wire is 280 cm long